how many liters of neon gas are in 7 moles of neon gas?


1
Expert's answer
2021-01-25T04:24:09-0500

At standard temperature pressure conditions, the volume of "n" moles of neon gas is:

"V= \frac{nRT}{p}" ,

where "R" is the ideal gas constant 8.314 J/(mol K), "T" is the temperature 273.15 K and "p" is the pressure 105 Pa. Therefore:

"V = \frac{7·8.314·273.15}{10^5} = 0.159" m3, or 1590 L.


Answer: at STP conditions, 7 moles of neon gas occupy 1590 L.